MAT 145: Topics In Contemporary Math
Trees
3) The restaurant from #2 has a higher-tier menu for $18.99. The diner still gets an
appetizer, main course, and desert, but there are more options for each, as shown below:
Appetizer
Shrimp Cocktail
Chicken Livers
French Onion Soup
Nachos
Main Course
16-oz sirloin
Small Prime Rib
Lasagna
Turkey Dinner
Shrimp Scampi
Fried Scallops
Dessert
Cheesecake
Pie a `la mode
Seasonal Fruit Parfait
How many different $18.99 specials are possible?

Answer by math_tutor2020(3817)


Given info:
Appetizer
Shrimp Cocktail
Chicken Livers
French Onion Soup
Nachos

Main Course
16-oz sirloin
Small Prime Rib
Lasagna
Turkey Dinner
Shrimp Scampi
Fried Scallops

Dessert
Cheesecake
Pie a `la mode
Seasonal Fruit Parfait

There are A = 4 appetizers, B = 6 main courses, and C = 3 desserts.

This means there are A*B*C = 4*6*3 = 24*3 = 72 specials priced at $18.99
If you were to make a tree diagram, then you'd have 72 different paths/branches.